AIPMT 2014 Q6. PLEASE PROVIDE A DETAILED SOLUTION....

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
@ anmol 
there are 2 set in 2014 pmt exam .,so, kindly mention which ques u are talking about . 
put an attachment always  while asking such type of questions . 
there is a baloon ques acc to me , i am solving that 
mg – F = ma (1st equation ) 
F- (m- m’ )g = (m-m’) a 
from (1) 
F-mg +m’g = ma – m’amg-ma-mg+m’g= ma-m’a 
m’(g+a) = 2ma 
so, m’ = 2ma/g+a 
so, option a turns out to be the correct option here . 
if u hv any other ques , post it with attachment .